Suppose that each team in a league has a demand curve for generic advertising (a league-wide nonteam–specific campaign) equal to Q=5,000-10p. If there are 40 teams in the league, and ads cost $325 each, how many ads will the teams want to purchase as a group? How would your answers change if league advertising were considered to be a private good?
Answer #1

Since advertisement is a public good which means there will be vertical summation

10P=5000-Q P=500-0.1Q Market demand P=20000-4Q and MC=325

Thus price=MC=325

P=325=20000-4Q Q=19675/4=4918.75 and Price=325

If it would have been a private good then market demand will be a horizontal summation of individual demand curve

Q=5000(40)-10(40)P=200000-400P

when P=325 then Q=200000-400(325)=70000